OpenStudy (aimmebgood):

Factor completely (a-3)^(2) - (4b)^(2)

OpenStudy (anonymous):

a lot like the last one it is the difference of two squares \[x^2-y^2=(x+y)(x-y)\] use \(x=(a-3)\) and \(y=2b\)

OpenStudy (theviper):

\((a-3)^{2} - (4b)^2\) \[=[(a-3)+(4b)][(a-3)-(4b)]\]

OpenStudy (anonymous):

ooops i made a mistake , it is \(4b)^2\) not \((4b^2\) so go with @TheViper answer

OpenStudy (theviper):

oh lol sir :)